A 12.5 V Battery operates s 14.2 ohms flashlight intermittently for a cumulative

time of 65.3 min. (a) What is the average current? (b) What quantity of electricity

is furnished by the battery?


1
Expert's answer
2022-03-28T14:24:41-0400

(1)We know that

"I=\frac{V}{R}=\frac{12.5}{14.2}=0.88A"

(2)

Power

"P=i^2R"

"P=\frac{12.5^2}{14.2}=11.003W"

Energy

"E=P\times t"


"E=11.003\times65.3\times60=43111.79J"